Suppose that domestic supply and demand for steelis given by the equations below, and the world price of steelis PWorld= $100. Demand: P = 400 –2QSupply: P = 40 + 4Qa.Graph domestic supply and demand.b.Solvefor the equilibrium price and quantity if the economy is closed to trade.c.If the country opens to trade, will it import or export steel? How much?d.Calculate the increase in total surplus (gains from trade) that will result from international trade in steel. e.In the domestic market, which group will likely be opposed to international trade? W
1
Expert's answer
2020-02-28T09:02:46-0500

a. The graph of domestic supply and demand is represented with the two lines, which intersects at the equilibrium point.

b. The equilibrium price and quantity without trade are:

400 - 2Q = 40 + 4Q,

6Q = 360,

Q = 60 units,

P = 400 - 2×60 = $280.

c.If the country opens to trade, then it will import steel.

At Pw = $100:

Qd = 200 - 100/2 = 150 units,

Qs = 100/4 - 10 = 15 units,

So, 150 - 15 = 135 units will be imported.

d. The increase in total surplus (gains from trade) that will result from international trade in steel is:

0.5×(150 - 15)×(280 - 100) = $12,150.

e. In the domestic market producers will likely be opposed to international trade.